The math team wants to buy new team t-shirts. Graphics by Andrew charges a $50 set-up fee and $8 per shirt. Team Andres charges a $40 set-up fee and $10 per shirt. How many shirts have to be ordered to make the costs the same from each company?
Answer by richwmiller(17219)

50+8x=40+10x
10=2x
5=x
5 shirts
check
50+40=40+50
ok